A 1000 kg truck is parked on a bridge 100 m long at a distance of 25 m from one end of
the bridge. The mass of the bridge is 200 kg m−1
. Calculate the reaction forces at the
supports at the two ends of the bridge. (Take g = 10 ms−2
)
1
Expert's answer
2019-10-28T11:24:28-0400

The force with which the truck acts on the bridge


"F=m\times g=1000\times10=10000 N"

Distributed load


"q=m\times g=200\times 10=2000 N"

Concentrate force


"Q=q\times l=2000\times100=200000 N"

The sum of the moments acting about one of the bridge supports (point a)


"∑M_A =0.25\times l\times F-l\times R_B+0.5\times l\times Q"

The reaction forces at the supports at the end of the bridge.


"R_B={(0.5\times l\times Q+0.25\times l\times F)}/{l}={(50\times 200000+25\times10000)}/{100}=102.5 kN"

The sum of the moments acting about one of the bridge supports (point b)

"∑M_B =0.75\times l\times F-l\times R_B+0.5\times l\times Q"

The reaction forces at the supports at the end of the bridge.


"R_A={(0.5\times l\times Q+0.75\times l\times F)}/{l}={(50\times 200000+75\times 10000)}/{100}=107.5 kN"